### v2

@@ 3.2.2 Equivalence to MINRES
The minimum residual (MINRES) method of Paige and Saunders 16 is a versatile approach for solving any linear systems/least-squares problems involving symmetric matrices. The iterates are constructed in a way that prevents unlucky breakdowns, guaranteeing the algorithm’s convergence to a normal solution. In fact, Facts 1 and 3.13 hold true for MINRES without the necessity to exclude any unfortunate situations, such as those specified in Algorithm 2.
@@
In essence, the above problem is solved using the QR factorization of $\tilde{\mathbf{T}}_{k}$. However, by exploiting its sparsity structure, we can simplify the computations and obtain an update procedure of the form $\mathbf{x}_{k}=\mathbf{x}_{k-1}+\tau_{k-1}\mathbf{d}_{k-2}$, for some step size $\tau_{k-1}$ and the update direction $\mathbf{d}_{k}$; see the works 16, 19 for more details.
For positive definite matrices, it is well-known that CR and MINRES, mathematically speaking, generate identical iterates 21. However, beyond PD settings, such an equivalence, at least to our knowledge, has not been thoroughly studied. In these general settings, the $\mathbf{A}$-weighted product .,.$\left\langle.,.\right\rangle_{\mathbf{A}}$ in CR is no longer a valid inner product and the equivalence between CR and MINRES is no longer automatic. In this section, we study the existence of such an equivalence in indefinite and/or inconsistent settings, and show that, under Algorithm 2, both CR and MINRES generate the same sequence of iterates for any symmetric systems.
To that end, we first demonstrate that, just like MINRES, subproblems of CR can also be formulated as 5 for any symmetric matrix $\mathbf{A}$. Let $\mathbf{P}_{k}=\begin{bmatrix}\mathbf{p}_{0}&\cdots&\mathbf{p}_{k-1}\end{bmatrix}$ and $\mathbf{R}_{k}=\begin{bmatrix}\mathbf{r}_{0}&\cdots&\mathbf{r}_{k-1}\end{bmatrix}$ be the matrices whose columns are, respectively, the update directions and the residual vectors from CR. Using Section 3-(a) and noting that $\mathbf{r}_{0}=\mathbf{b}$, we can write $\mathbf{x}=\mathbf{P}_{k}\mathbf{z}$ for $\mathbf{z}\in\mathbb{R}^{k}$ and
@@
$\mathbf{z}_{k}$ $\triangleq\underset{\mathbf{z}\in\mathbb{R}^{k}}{\text{argmin}}\frac{1}{2}\left\langle\mathbf{z},\mathbf{P}_{k}^{{\top}}\mathbf{A}^{2}\mathbf{P}_{k}\mathbf{z}\right\rangle-\left\langle\mathbf{z},\mathbf{P}_{k}^{{\top}}\mathbf{A}\mathbf{R}_{k}\mathbf{e}_{1}\right\rangle.$ (8)
By Section 3-(c), the matrix $\mathbf{D}_{k}\triangleq\mathbf{P}_{k}^{{\top}}\mathbf{A}^{2}\mathbf{P}_{k}$ is diagonal, i.e.,
@@
Note that, by Lemma 3.11, the matrix $\mathbf{D}_{k}$ is always invertible prior to termination. Also, from Properties 3-(a) and 3-(c), it follows that the matrix $\mathbf{K}_{k}\triangleq\mathbf{P}_{k}^{{\top}}\mathbf{A}\mathbf{R}_{k}$ is lower triangular, i.e.,
@@
Lemma 3.19.
Let $\mathbf{b}\in\text{Range}(\mathbf{A})$. Then, under Algorithm 2, CR and MINRES generate the same sequence of iterates.
Proof 3.20.
@@
$\underset{\mathbf{y}\in\mathbb{R}^{k}}{\text{argmin}}\frac{1}{2}\left\|\mathbf{b}-\mathbf{AW}_{k}\mathbf{y}\right\|^{2}=\underset{\mathbf{y}\in\mathbb{R}^{k}}{\text{argmin}}\frac{1}{2}\left\langle\mathbf{y},\mathbf{W}_{k}^{{\top}}\mathbf{A}^{2}\mathbf{W}_{k}\mathbf{y}\right\rangle-\left\langle\mathbf{y},\mathbf{W}_{k}^{{\top}}\mathbf{Ab}\right\rangle.$
By the assumption on $\mathbf{b}$, $\text{Span}(\mathbf{W}_{k})\subseteq\text{Range}(\mathbf{A})$, which implies $\mathbf{W}_{k}^{{\top}}\mathbf{A}^{2}\mathbf{W}_{k}\succ\mathbf{0}$. Hence, the above problem admits a unique solution $\mathbf{y}_{k}$, resulting in a unique $\mathbf{x}_{k}$.
Hence, when $\mathbf{b}\in\text{Range}(\mathbf{A})$, there can be only one unique minimizer of 5 in $\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$ for $1\leq k\leq g$. However, the case for inconsistent systems may appear less straightforward. Specifically, when $\mathbf{b}\not\in\text{Range}(\mathbf{A})$, there may be more than one solution to 5. The crucial step, now, is to show that, despite possibly many solutions to 5, the minimizer from CR is indeed the same as the one from MINRES, i.e., the sequence of iterates $\{\mathbf{x}_{k-1}\}_{k=1}^{k=g}$ for both algorithms, in exact arithmetic, coincide. To do this, we first establish a counterpart of Lemma 3.16 for MINRES (Lemma 3.21) and then leverage Lemmas 3.19, 3.16 and 3.21 to establish our claim (Theorem 3.23). A diagram depicting our proof strategy is given in Figure 1.
@@
Figure 1: Our proof strategy for establishing equivalence between MINRES and CR. To prove Theorem 3.23, we use Lemmas 3.19, 3.16 and 3.21.
Lemma 3.21.
@@
Theorem 3.23.
Under Algorithm 2, CR and MINRES generate the same sequence of iterates.
Proof 3.24.
@@
Using Lemmas 3.19, 3.16 and 3.21, we have $\mathbf{A}^{\dagger}\mathbf{A}\mathbf{x}_{k-1}^{\text{MINRES}}=\mathbf{A}^{\dagger}\mathbf{A}\mathbf{x}_{k-1}^{\text{CR}}$, which implies
@@
Now, from linear independence of $\{\mathbf{A}^{i-1}\tilde{\mathbf{b}}\}_{i=1}^{k-1}$, we get $a_{i-1}=b_{i-1}$ for $11\leq i\leq k-$ 1.
Theorem 3.23 implies that all properties established for MINRES in prior works, such as non-positive curvature detection, monotonicity, and lifting procedures in 34, 19, are also readily applicable to CR.
After establishing the equivalence between MINRES and CR under Algorithm 2, it is of theoretical interest to investigate the behavior of MINRES when CR experiences an unlucky breakdown. To address this, we present the following two lemmas that characterize this behavior.
Lemma 3.25.
@@
where the first equality follows from Section 3-(c). By Lemma 3.25, as long as $\left\langle\mathbf{r}_{i},\mathbf{Ar}_{i}\right\rangle\neq 0$ for $20\leq i\leq k-$ 2, the columns of $\mathbf{R}_{k}$ are linearly independent and $\mathbf{Z}_{k}$ must be full rank. Hence, $\left\langle\mathbf{r}_{k-1},\mathbf{Ar}_{k-1}\right\rangle=0$ if and only if $\mathbf{T}_{k}$ is singular.
@@ 4 Numerical Experiments
In this section, we conduct a series of experiments to verify aspects of the theory developed in this paper and to explore the empirical performance of the methods discussed in various settings. In Section 4.1, using synthetic examples, our aim is to empirically verify Lemmas 3.3 and 3.1. Additionally, we will investigate the stability of these methods and identify instances where the theoretical predictions deviate significantly from practical observations. Moving to Section 4.2, our focus remains on the same objectives, but within the context of real-world applications. It is crucial to note that our primary goal here is to observe the behavior of these methods in their original form, without incorporating any additional enhancements such as reorthogonalization strategies (if applicable).
Algorithms are terminated either upon reaching a maximum number of iterations or when a specific termination condition is met. For CG, this condition is chosen as $\min\{\|\mathbf{r}_{k}\|/\|\mathbf{b}\|,\|\mathbf{Ap}_{k}\|/\|\mathbf{Ab}\|\}$ (cf. Facts 1 and 3.3), and for CR/MINRES, it is $\min\{\|\mathbf{r}_{k}\|/\|\mathbf{b}\|,\|\mathbf{Ar}_{k}\|/\|\mathbf{Ab}\|\}$ (cf. Facts 1 and 3.13).
@@ 4.1 Synthetic Problems
In this section, we experiment with a series of simple and synthetic diagonal matrices to verify Lemmas 3.3 and 3.1 and to explore the stability of these methods in several instances. To validate Lemmas 3.3 and 3.1, we focus on small-size problems in Section 4.1.1 to avoid any unwanted numerical issues. Subsequently, we extend our analysis to larger-size matrices in Sections 4.1.3 and 4.1.2 to investigate the stability of these methods across a substantial number of iterations.
@@
4.1.1 CG at Termination: Validating Lemmas 3.3 and 3.1
In this section, we validate Lemmas 3.3 and 3.1. For this, we consider $d=10$, and generate six matrices by randomly zeroing out $m=5,4,3,2,1$ and 00 diagonal elements of $\mathbf{A}$. The results are depicted in Figures 2 and 3 where we consider PSD and indefinite settings, respectively. In all these examples, as predicted by Lemma 3.3, $\|\mathbf{Ap}_{k}\|$ indeed vanishes at termination, while $\|\mathbf{Ar}_{k}\|$ remains bounded away from zero in all but consistent case, validating Theorem 3.1.
Figure 2: Experiments for Section 4.1.1 to verify Lemmas 3.3 and 3.1. The matrices are PSD with $d=10$. From the top left to the bottom right, the grades are 6, 7, 8, 9, 10 and 10. The bottom right is a positive definite matrix.
Figure 3: Experiments for Section 4.1.1 to verify Lemmas 3.3 and 3.1. The matrices are indefinite with $d=10$. From the top left to the bottom right, the grades are 6, 7, 8, 9, 10 and 10. The bottom right is a full rank indefinite matrix.
@@ 4.1.2 Numerical Instability of CG
In this section, we consider larger size matrices than those in Section 4.1.1 to explore the behavior of CG across a larger number of iterations. Specifically, we set $d=100$ and generate a series of matrices by zeroing out $m=90,80,60,20,1$ and 00 diagonal entries. The maximum number of iterations allowed is capped at $g$, i.e., the corresponding grade for each example, and the tolerance for the termination condition $\min\{\|\mathbf{r}_{k}\|/\|\mathbf{b}\|,\|\mathbf{Ap}_{k}\|/\|\mathbf{Ab}\|\}$ is $10^{-8}$. The results are gathered in Figures 4, 5 and 6, corresponding to PSD, indefinite, and slightly indefinite settings, respectively. In all these examples, when $g$ is small, the theoretical prediction of Lemma 3.3 aligns with the observed behavior. However, as $g$ increases, necessitating more iterations from CG, a clear discrepancy emerges. Unless the system is consistent, the implications outlined in Lemma 3.3 do not accurately reflect practical outcomes, highlighting the presence of substantial numerical inaccuracies within CG iterations. Nonetheless, an interesting observation from Figures 6 and 5 is that the presence of negative eigenvalues appears to induce a form of oscillatory behavior in CG. While the depicted quantities remain bounded away from zero, they do not exhibit the pronounced growth observed in Figure 4.

### v4

@@ 3.2.2 Equivalence to MINRES
The minimum residual (MINRES) method of Paige and Saunders [34] is a versatile approach for solving any linear systems/least-squares problems involving symmetric matrices. The iterates are constructed in a way that prevents unlucky breakdowns, guaranteeing the algorithm’s convergence to a normal solution. In fact, Facts 1 and 3 hold true for MINRES without the necessity to exclude any unfortunate situations, such as those specified in Assumption 1.
@@
In essence, the above problem is solved using the QR factorization of $\tilde{\mathbf{T}}_{k}$. However, by exploiting its sparsity structure, we can simplify the computations and obtain an update procedure of the form $\mathbf{x}_{k}=\mathbf{x}_{k-1}+\tau_{k-1}\mathbf{d}_{k-2}$, for some step size $\tau_{k-1}$ and the update direction $\mathbf{d}_{k}$; see the works [34, 28] for more details.
For positive definite matrices, it is well-known that CR and MINRES, mathematically speaking, generate identical iterates [14]. However, beyond PD settings, such an equivalence, at least to our knowledge, has not been thoroughly studied. In these general settings, the $\mathbf{A}$-weighted product .,.$\left\langle.,.\right\rangle_{\mathbf{A}}$ in CR is no longer a valid inner product and the equivalence between CR and MINRES is no longer automatic. In this section, we study the existence of such an equivalence in indefinite and/or inconsistent settings, and show that, under Assumption 1, both CR and MINRES generate the same sequence of iterates for any symmetric systems.
To that end, we first demonstrate that, just like MINRES, subproblems of CR can also be formulated as 5 for any symmetric matrix $\mathbf{A}$. Let $\mathbf{P}_{k}=\begin{bmatrix}\mathbf{p}_{0}&\cdots&\mathbf{p}_{k-1}\end{bmatrix}$ and $\mathbf{R}_{k}=\begin{bmatrix}\mathbf{r}_{0}&\cdots&\mathbf{r}_{k-1}\end{bmatrix}$ be the matrices whose columns are, respectively, the update directions and the residual vectors from CR. Using Property 1-(a) and noting that $\mathbf{r}_{0}=\mathbf{b}$, we can write $\mathbf{x}=\mathbf{P}_{k}\mathbf{z}$ for $\mathbf{z}\in\mathbb{R}^{k}$ and
@@
$\mathbf{z}_{k}$ $\triangleq\underset{\mathbf{z}\in\mathbb{R}^{k}}{\text{argmin}}\frac{1}{2}\left\langle\mathbf{z},\mathbf{P}_{k}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\mathbf{A}^{2}\mathbf{P}_{k}\mathbf{z}\right\rangle-\left\langle\mathbf{z},\mathbf{P}_{k}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\mathbf{A}\mathbf{R}_{k}\mathbf{e}_{1}\right\rangle.$ (8)
By Property 1-(c), the matrix $\mathbf{D}_{k}\triangleq\mathbf{P}_{k}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\mathbf{A}^{2}\mathbf{P}_{k}$ is diagonal, i.e.,
@@
Note that, by Lemma 3, the matrix $\mathbf{D}_{k}$ is always invertible prior to termination. Also, from Properties 1-(a) and 1-(c), it follows that the matrix $\mathbf{K}_{k}\triangleq\mathbf{P}_{k}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\mathbf{A}\mathbf{R}_{k}$ is lower triangular, i.e.,
@@
Lemma 5.
Let $\mathbf{b}\in\text{Range}(\mathbf{A})$. Then, under Assumption 1, CR and MINRES generate the same sequence of iterates.
Proof.
@@
$\underset{\mathbf{y}\in\mathbb{R}^{k}}{\text{argmin}}\frac{1}{2}\left\|\mathbf{b}-\mathbf{AW}_{k}\mathbf{y}\right\|^{2}=\underset{\mathbf{y}\in\mathbb{R}^{k}}{\text{argmin}}\frac{1}{2}\left\langle\mathbf{y},\mathbf{W}_{k}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\mathbf{A}^{2}\mathbf{W}_{k}\mathbf{y}\right\rangle-\left\langle\mathbf{y},\mathbf{W}_{k}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\mathbf{Ab}\right\rangle.$
By the assumption on $\mathbf{b}$, $\text{Span}(\mathbf{W}_{k})\subseteq\text{Range}(\mathbf{A})$, which implies $\mathbf{W}_{k}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\mathbf{A}^{2}\mathbf{W}_{k}\succ\mathbf{0}$. Hence, the above problem admits a unique solution $\mathbf{y}_{k}$, resulting in a unique $\mathbf{x}_{k}$. ∎
Hence, when $\mathbf{b}\in\text{Range}(\mathbf{A})$, there can be only one unique minimizer of 5 in $\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$ for $1\leq k\leq g$. However, the case for inconsistent systems may appear less straightforward. Specifically, when $\mathbf{b}\not\in\text{Range}(\mathbf{A})$, there may be more than one solution to 5. The crucial step, now, is to show that, despite possibly many solutions to 5, the minimizer from CR is indeed the same as the one from MINRES, i.e., the sequence of iterates $\{\mathbf{x}_{k-1}\}_{k=1}^{k=g}$ for both algorithms, in exact arithmetic, coincide. To do this, we first establish a counterpart of Lemma 4 for MINRES (Lemma 6) and then leverage Lemmas 5, 4 and 6 to establish our claim (Theorem 5). A diagram depicting our proof strategy is given in Figure 1.
@@
Figure 1: Our proof strategy for establishing equivalence between MINRES and CR. To prove Theorem 5, we use Lemmas 5, 4 and 6.
Lemma 6.
@@
Theorem 5.
Under Assumption 1, CR and MINRES generate the same sequence of iterates.
Proof.
@@
Using Lemmas 5, 4 and 6, we have $\mathbf{A}^{\dagger}\mathbf{A}\mathbf{x}_{k-1}^{\text{MINRES}}=\mathbf{A}^{\dagger}\mathbf{A}\mathbf{x}_{k-1}^{\text{CR}}$, which implies
@@
Now, from linear independence of $\{\mathbf{A}^{i-1}\tilde{\mathbf{b}}\}_{i=1}^{k-1}$, we get $a_{i-1}=b_{i-1}$ for $11\leq i\leq k-$ 1. ∎
Theorem 5 implies that all properties established for MINRES in prior works, such as non-positive curvature detection, monotonicity, and lifting procedures in [28, 26], are also readily applicable to CR.
After establishing the equivalence between MINRES and CR under Assumption 1, it is of theoretical interest to investigate the behavior of MINRES when CR experiences an unlucky breakdown. To address this, we present the following two lemmas that characterize this behavior.
Lemma 7.
@@
where the first equality follows from Property 1-(c). By Lemma 7, as long as $\left\langle\mathbf{r}_{i},\mathbf{Ar}_{i}\right\rangle\neq 0$ for $20\leq i\leq k-$ 2, the columns of $\mathbf{R}_{k}$ are linearly independent and $\mathbf{Z}_{k}$ must be full rank. Hence, $\left\langle\mathbf{r}_{k-1},\mathbf{Ar}_{k-1}\right\rangle=0$ if and only if $\mathbf{T}_{k}$ is singular. ∎
@@ 4 Numerical Experiments
In this section, we conduct a series of experiments to verify aspects of the theory developed in this paper and to explore the empirical performance of the methods discussed in various settings. While numerical inaccuracies due to finite precision arithmetic in iterative methods have long been well known and thoroughly studied, our main focus is to highlight the additional and unexpectedly catastrophic instability of CG in inconsistent settings–behavior not typically observed in consistent settings–and compare it to the CR alternative. In particular, we set out to demonstrate the following:
  1. 1.
    CR is generally more stable than CG in most cases, regardless of consistency.
  2. 2.
    Unlike CR, CG exhibits extreme instability under inconsistent systems.
  3. 3.
    Unlike CR, CG can be unstable even for PD matrices with a few small eigenvalues.
In Section 4.1, using synthetic examples, our aim is to empirically verify Lemmas 1 and 1. Additionally, we will investigate the stability of these methods and identify instances where the theoretical predictions deviate significantly from practical observations. Moving to Section 4.2, our focus remains on the same objectives, but within the context of real-world applications. It is crucial to note that our primary goal here is to observe the behavior of these methods in their original form, without incorporating any additional enhancements such as reorthogonalization strategies (if applicable).
Algorithms are terminated either upon reaching a maximum number of iterations or when a specific termination condition is met. For CG, this condition is chosen as $\min\{\|\mathbf{r}_{k}\|/\|\mathbf{b}\|,\|\mathbf{Ap}_{k}\|/\|\mathbf{Ab}\|\}$ (cf. Facts 1 and 1), and for CR/MINRES, it is $\min\{\|\mathbf{r}_{k}\|/\|\mathbf{b}\|,\|\mathbf{Ar}_{k}\|/\|\mathbf{Ab}\|\}$ (cf. Facts 1 and 3).
@@ 4.1 Synthetic Problems
In this section, we experiment with a series of simple and synthetic diagonal matrices to verify Lemmas 1 and 1 and to explore the stability of these methods in several instances. To validate Lemmas 1 and 1, we focus on small-size problems in Section 4.1.1 to avoid any unwanted numerical issues. Subsequently, we extend our analysis to larger-size matrices in Sections 4.1.3 and 4.1.2 to investigate the stability of these methods across a substantial number of iterations.
@@
4.1.1 CG at Termination: Validating Lemmas 1 and 1
In this section, we validate Lemmas 1 and 1. For this, we consider $d=10$, and generate six matrices by randomly zeroing out $m=5,4,3,2,1$ and 00 diagonal elements of $\mathbf{A}$. The results are depicted in Figures 2 and 3 where we consider PSD and indefinite settings, respectively. In all these examples, as predicted by Lemma 1, $\|\mathbf{Ap}_{k}\|$ indeed vanishes at termination, while $\|\mathbf{Ar}_{k}\|$ remains bounded away from zero in all but consistent case, validating Theorem 1.
Figure 2: Experiments for Section 4.1.1 to verify Lemmas 1 and 1. The matrices are PSD with $d=10$. From the top left to the bottom right, the grades are 6, 7, 8, 9, 10 and 10. The bottom right is a positive definite matrix.
Figure 3: Experiments for Section 4.1.1 to verify Lemmas 1 and 1. The matrices are indefinite with $d=10$. From the top left to the bottom right, the grades are 6, 7, 8, 9, 10 and 10. The bottom right is a full rank indefinite matrix.
@@ 4.1.2 Numerical Instability of CG
In this section, we consider larger size matrices than those in Section 4.1.1 to explore the behavior of CG across a larger number of iterations. Specifically, we set $d=100$ and generate a series of matrices by zeroing out $m=90,80,60,20,1$ and 00 diagonal entries. The maximum number of iterations allowed is capped at $g$, i.e., the corresponding grade for each example, and the tolerance for the termination condition $\min\{\|\mathbf{r}_{k}\|/\|\mathbf{b}\|,\|\mathbf{Ap}_{k}\|/\|\mathbf{Ab}\|\}$ is $10^{-8}$. The results are gathered in Figures 4, 5 and 6, corresponding to PSD, indefinite, and slightly indefinite settings, respectively. In all these examples, when $g$ is small, the theoretical prediction of Lemma 1 aligns with the observed behavior. However, as $g$ increases, necessitating more iterations from CG, a clear discrepancy emerges. Unless the system is consistent, the implications outlined in Lemma 1 do not accurately reflect practical outcomes, highlighting the presence of substantial numerical inaccuracies within CG iterations. Nonetheless, an interesting observation from Figures 6 and 5 is that the presence of negative eigenvalues appears to induce a form of oscillatory behavior in CG. While the depicted quantities remain bounded away from zero, they do not exhibit the pronounced growth observed in Figure 4.